Using statkeys create a 95% confidence interval estimate for the true proportion of successful first attempt intubations in obese patients. Dataset:attempt1_S_F: 1 1 0 0 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 0 1 1 1 1 1 1 1 1 1 0 1 1 1 1 1 1 1 0 1 1 1 1 1 0 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 0 1 1 1 1 1 1 1 1 0 1 0 1 1 1 1 1 1 0 0 1
Answer #1

X = 88

n = 99

p^ = X/n = 88/99 = 0.8889

95% confidence interval

p^ +- 1.96* sqrt(p^ * q^/n)

= (8/9 +- 1.96* sqrt(8/9*1/9/99)

= (0.8889 +- 0.0619)

= (0.8270,0.9508)

Please give me a thumbs-up if this helps you out. Thank you! :)